One number is 3 times another. Their product is 3 times their sum. Find the numbers x and y. (Assume x > y.)
Answer by CubeyThePenguin(3113)

x = 3y
xy = 3(x + y)

Substitute x = 3y into the second equation.

3y(y) = 3(3y + y)
4y^2 = 3(4y)
4y^2 = 12y
y^2 - 3y = 0
y(y - 3) = 0

(x, y) = (0, 0) or (3, 9)